OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned? 44 36 76 60 There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $390. How many sophomores bought tickets? 100 90 60 30

OpenStudy (anonymous):

m = number of men w = number of women m + w = 100 2m - 3w = 20 m = 100 - w Put this equation into the m part of the second equation 2(100 - w) - 3w = 20 200 - 2w - 3w = 20 200 - 20 = 5w 180 = 5w

OpenStudy (anonymous):

When you divide 180 by 5 you get your answer

OpenStudy (anonymous):

Oh this is for the first one

OpenStudy (anonymous):

36 wht about second one

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

f + s = 150 2f + 3s = 390 f = 150 - s 2f + 3s = 390 2(150 - s) + 3s = 390 2(150 - s) + 3s = 390 300 - 2s + 3s = 390 300 + s = 390

OpenStudy (anonymous):

Subtract 300 from 390

OpenStudy (anonymous):

equals 90

OpenStudy (anonymous):

Yay ! Thank s!